# AI-Based Beam-Level and Cell-Level Mobility Management for High Speed Railway Communications

### v1

@@ Abstract
High-speed railway (HSR) communications are pivotal for ensuring rail safety, operations, maintenance, and delivering passenger information services. The high speed of trains creates rapidly time-varying wireless channels, increases the signaling overhead, and reduces the system throughput, making it difficult to meet the growing and stringent needs of HSR applications. In this article, we explore artificial intelligence (AI)-based beam-level and cell-level mobility management suitable for HSR communications, including the use cases, inputs, outputs, and key performance indicators (KPI)s of AI models. Particularly, in comparison to traditional down-sampling spatial beam measurements, we show that the compressed spatial multi-beam measurements via compressive sensing lead to improved spatial-temporal beam prediction. Moreover, we demonstrate the performance gains of AI-assisted cell handover over traditional mobile handover mechanisms. In addition, we observe that the proposed approaches to reduce the measurement overhead achieve comparable radio link failure performance with the traditional approach that requires all the beam measurements of all cells, while the former methods can save 50% beam measurement overhead.
@@
Artificial intelligence, beam-level mobility management, cell-level mobility management.I Introduction
@@
High speed railway (HSR) provides people with safe, fast, comfortable and economical modes of transportation, and has been widely developed around the world. Nevertheless, mobile communication systems are pivotal for ensuring rail safety, operations, maintenance, and delivering passenger information services. Emerging applications, such as railway multimedia dispatching, high-definition video monitoring on board and trackside, have posed stricter requirements on the reliability, data rate, and spectrum efficiency of the next-generation, dedicated mobile communication systems for HSR. However, the high speed of trains entails formidable challenges due to the time-varying nature of wireless channels in HSR communications. These challenges include the increased signaling overhead and reduced system throughput.
@@
Thanks to the abundant spectrum available at higher frequencies, communication systems using the millimeter wave (mmWave) technology avail of faster data transmission speeds and reduced latency compared to traditional low-frequency communications. For HSR communications, mmWave technology underpins increased data transmission rates that meet the demands of onboard applications and services, and provides low latency that is indispensable for real-time communication requirements. However, mmWave communication systems suffer from significant path loss, which restricts the effective range of transmission. To tackle this problem, beamforming with large-scale antenna arrays can be employed to generate finely directional beams, mitigating propagation loss and expanding signal coverage. However, the high mobility of trains entails pressing challenges in mobility management for HSR communications. While being connected, a train has to address two types of mobility, i.e., the cell-level and beam-level mobility as specified in the 3rd Generation Partnership Project (3GPP) Release 17 for 5G systems[1]. Beam-level mobility involves dynamically switching beams in which the train establishes a link path to a new beam within the same cell, ensuring communication with high performance. On the other hand, cell-level mobility focuses on cell handover between different cells. As trains move from one coverage zone to another, it is crucial to have a smooth cell handover to sustain continuous connectivity experience.
@@
The management of mobility at the beam and cell level in 5G is based on the use of periodic channel state information reference signals (CSI-RS) and synchronization signal blocks (SSBs). Unfortunately, the resulting large latency and high overhead are prohibitive for HSR communications, and, hence, innovative approaches are urgently needed to address these limitations. Deep neural networks (DNNs) have found extensive applications in areas, such as computer vision and natural language processing, owing to their exceptional ability to extract features from intricate and high-dimensional data. In recent years, DNNs have also proven their effectiveness in capturing the complex and nonlinear characteristics of wireless channels to facilitate beam-level mobility management. For example, a multi-layer long short term memory (LSTM) network was utilized to predict the optimal beam change for future instances [2]. Studies on advanced artificial intelligence (AI)-based cell-level mobility management are relatively few. For example, [3] utilized convolutional neural network (CNN)-based auto-encoders for dimension reduction and LSTM networks to predict the appropriate base station (BS) for cell handover. Recently, the 3GPP 5G NR organization has launched studies on AI-based cell-level mobility management for highly reliable handovers in Release 19. However, these AI-based mobility management methods fail to exploit the specific characteristics in HSR communications. For instance, trains typically maintain stable high-speed motion on fixed tracks and have different speed levels to meet different needs. Mobility management methods for HSR communications are expected to leverage case-specific characteristics in the design phase or, alternatively, learn from data to yield improved performance in a data-driven manner.
In this article, we present the latest progress in AI-based beam-level and cell-level mobility management suitable for HSR communications. In Section II, we reveal the potential of AI-based spatial-temporal beam prediction for HSR communications. Particularly, in comparison to the traditional down-sampling spatial beam measurements, we show that the compressed spatial multi-beam measurements via compressive sensing (CS) can lead to improved spatial-temporal beam prediction in HSR communications. In Section III, we investigate AI-based cell-level mobility management for HSR scenarios, including use cases, the input/output of the AI model, evaluation indicators, and enhanced schemes. Particularly, we demonstrate the performance gains of AI-assisted cell handover over traditional cell handover mechanisms for HSR scenarios. Section V introduces the challenges and opportunities encountered in AI-based mobility management, followed by our conclusions in Section VI.
II AI-Based Beam-Level Mobility Management for HSR Communications
II-A Use Cases in HSR Communications
Beam-level mobility enables a user equipment (UE) to communicate with a BS via the optimal beam pair, which involves low-level beam operations rather than handover. In HSR communications, with some coarse-grained beam quality measurements of the serving cell, AI can be utilized to estimate the finely-grained optimal beam pair in the current instance or predict the optimal beam pair in future instances, which leads to reduced measurement overhead and enhanced throughput.
There are two primary use cases for AI models applied in beam-level mobility management, i.e., temporal beam prediction and spatial beam prediction. For temporal beam prediction, where the beam quality information of previous instances is utilized to predict the quality of beams in one or more future instances, most AI-based methods employ recurrent neural networks (RNNs) and their variants, such as LSTM and gated recurrent unit (GRU) models. This is driven by the ability of RNN models to identify temporal patterns in time series data, allowing them to gather extensive information. A two-layer convolutional LSTM network can be used for predicting the beam quality across spatial and temporal dimensions as detailed in [4]. The combination of a CNN and LSTM network was proposed in [5] for predicting the optimal beam ID across temporal dimensions. A multi-layer LSTM network was applied in [6] for forecasting the optimal future beam. Spatial beam prediction exploits spatial correlations among adjacent beams to estimate the quality of all beams using incomplete beam data. CNNs and networks with fully connected (FC) layers have been commonly employed for this purpose [7, 8, 9]. Due to the high mobility in HSR communications, the extensive signaling overhead and the time required for frequent beam sweeping become prohibitive. For these reasons, temporal beam prediction offers greater benefits in terms of resource conservation and stable performance in HSR communications.
II-B Inputs and Outputs in HSR Communications
Different tasks in beam-level mobility management require distinct input/output into/of the AI models in HSR communications. According to the ongoing discussions within 3GPP, there are two vital concepts for beam management, namely Set A and Set B. Set A comprises all narrow beams from the predefined codebook, while Set B encompasses the beams that have been measured and are utilized as inputs into the AI models. Various alternative methods exist with respect to the relationship between Set A and Set B. For instance, the AI model could receive the input of reference signal received power (RSRP) measurements of partial narrow beams, making Set B as a subset of Set A. This is the most commonly used method in HSR communications, which greatly reduces the measurement overhead. Another option is to utilize RSRP measurements of wide beams to estimate the quality of narrow beams, where Set B and Set A are distinct. In addition, Set B can be the same as Set A, whenever the RSRP measurements of all narrow beams are the inputs into the AI models.
Note that the first two types of inputs are employed in temporal and spatial beam prediction models, while the last is only for temporal beam prediction. Instead of using RSRP, the complex received signal can be used as input into the AI model, which will be discussed in the CS-based method in the next subsection. In addition, supplementary auxiliary information can be incorporated to improve the prediction accuracy for HSR communications. In [10], the position of the UE, its mobility vector, the signal-to-noise ratio (SNR), and the beam index currently in use were utilized to forecast the state of the link connection. Furthermore, measurements from light detection and range (LIDAR) and various sensors, images of the environment taken by cameras, and information from low-frequency channels could also be employed as extra inputs into the AI model to enhance the prediction accuracy.
There are two types of outputs in the AI model design for beam-level mobility management. The most common outputs are weights associated with different beams, where a higher weight suggests a better beam. Additionally, line-of-sight (LoS) is the primary mode of signal propagation in HSR communications and beam-selection is easier in the LoS setting. Therefore, the AI model can be utilized to output the estimated state of each path, i.e., LoS or non-LoS, to enhance the beam prediction [11, 9].
II-C AI Enhancement for HSR Communications
@@
Beam-level mobility management enhancement involves three types of beam search methods: traditional methods including exhaustive search and hierarchical search, advanced numerical methods, and AI-based methods. The exhaustive beam search method examines all the candidate beams in the predefined codebook one by one to obtain beam quality information. The hierarchical beam search method provides a more efficient alternative, which first searches within a wide beam range and then further searches for a narrow beam within the angular range of the selected wide beam. However, traditional beam search algorithms suffer from a large training overhead, which is not suitable for HSR communications which require frequent beam switching. In addition, enhancement techniques, based on numerical methods have been proposed. More specifically, [12] exploited the tree search method for efficient beam training based on the hierarchical codebook, whilst [13] proposed an optimal beam training model based on Markov’s decision process. These numerical approaches mainly target general communication environments, making it challenging to tailor them for particular communication scenarios.
@@
Here, we investigate DL-based spatial-temporal beam prediction using compressed multi-beam channel measurements. DL is employed to output the beam quality in the future slots using historical multi-beam channel measurements in the compressed domain, where the CS measurement matrix can be learned jointly. In the training phase, the complete beam channel information from historical instances serves as input into the model. The CS matrix serves as the first layer, namely the linear compression layer. The compressed channel information is fed into the remaining neural network to reconstruct the complete RSRP information of all candidate beams for a future instance. As a learnable parameter, the CS measurement matrix is jointly optimized during the training of the model. In the inference phase, the trained measurement matrices are applied at the transceivers and the compressed measurements are fed into the trained neural network to conduct beam prediction. Compared with the traditional sweeping method, the multi-beam compression offers to the beam prediction network richer information with the same number of beam measurements. The structure of the proposed AI model for beam-level mobility management, namely CSAI, is illustrated in Fig. 2(d).
@@
Now, we compare the performance of AI-based spatial-temporal beam prediction methods, whose structures are shown in in Fig. 2. The AI-based methods in Fig. 2(a) [6] and Fig. 2(b) [7] employ the traditional LSTM and CNN architecture, respectively, to predict the high-resolution beams. Compared with the CSAI in Fig. 2(d), the AI-based method in Fig. 2(c) [4] has the same network structure except for the beam measurement module. In the experiments, we consider the UMa scenario for outdoor 30 GHz mmWave communications, while the UE follows a straight-line trajectory which aligns with the HSR scenario. Four distinct datasets, representing UE speeds of 60km/h, 120km/h, 350km/h and 500km/h, are utilized in the evaluation. These datasets comprise 40,000 training samples, 5,000 validation samples, and 5,000 test samples. The batch size is set to 256, and the training runs for a maximum of 800 epochs. The learning rate changes dynamically from to , utilizing the adaptive moment estimation optimizer. We consider the prediction accuracy for the best beam with a compression ratio111The compression ratio is defined as the ratio of the number of input beam measurements to the number of all beam measurements. of . The simple non-AI approach straightforwardly uses the best one of the partially measured present beams as the prediction. As shown in Fig. 3, the AI-based spatial-temporal beam prediction methods outperform significantly the traditional non-AI method. As the speed of the movement increases, all methods generally show an anticipated decrease in the prediction accuracy. The CSAI demonstrates superior performance for all the cases. At the speed of 500km/h, the CSAI is still able to yield a prediction accuracy of 83.76%.
III AI-Based Cell-level Mobility Management in HSR Communications
III-A Use Cases in HSR Communications

### v2

# AI-Driven Mobility Management for High-Speed Railway Communications: Compressed Measurements and Proactive Handover

@@ Abstract
High-speed railway (HSR) communications are pivotal for ensuring rail safety, operations, maintenance, and delivering passenger information services. The high speed of trains creates rapidly time-varying wireless channels, increases the signaling overhead, and reduces the system throughput, making it difficult to meet the growing and stringent needs of HSR applications. In this article, we explore artificial intelligence (AI)-based beam-level and cell-level mobility management suitable for HSR communications. Particularly, we propose a compressed spatial multi-beam measurements scheme via compressive sensing for beam-level mobility management in HSR communications. In comparison to traditional down-sampling spatial beam measurements, this method leads to improved spatial-temporal beam prediction accuracy with the same measurement overhead. Moreover, we propose a novel AI-based proactive handover scheme to predict handover events and reduce radio link failure (RLF) rates in HSR communications. Compared with the traditional event A3-based handover mechanism, the proposed approach significantly reduces the RLF rates which saves 50% beam measurement overhead.
@@
Artificial intelligence, compressive sensing, mobility management.I Introduction
@@
High-speed railway (HSR) provides people with safe, fast, comfortable and economical modes of transportation, and has been widely developed around the world. Nevertheless, mobile communication systems are pivotal for ensuring rail safety, operations, maintenance, and delivering passenger information services. Emerging applications, such as railway multimedia dispatching and high-definition video monitoring on board and trackside, have posed stricter requirements on the reliability, data rate, and spectral efficiency of the next-generation, dedicated mobile communication systems for HSR. However, the high speed of trains entails formidable challenges due to the time-varying nature of wireless channels in HSR communications. These challenges include the increased signaling overhead and reduced system throughput.
@@
Thanks to the abundant spectrum available at higher frequencies, communication systems using the millimeter wave (mmWave) technology avail of faster data transmission speeds and reduced latency compared to traditional low-frequency communications. For HSR communications, mmWave frequency bands can underpin increased data transmission rates that meet the demands of onboard applications and services, and provides low latency that is indispensable for real-time communication requirements. However, mmWave communication systems suffer from significant path loss, which restricts the effective range of transmission. To tackle this problem, beamforming with large-scale antenna arrays can be employed to generate finely directional beams, mitigating propagation loss and expanding signal coverage. However, the high mobility of trains entails pressing challenges in mobility management for HSR communications. While being connected, a train has to address two types of mobility, i.e., the beam-level and cell-level mobility as specified in the 3rd Generation Partnership Project (3GPP) Release 17 for fifth generation (5G) systems[1]. Beam-level mobility involves dynamically switching beams in which the train establishes a link path to a new beam within the same cell, ensuring communication with high performance. On the other hand, cell-level mobility focuses on cell handover between different cells. As trains move from one coverage zone to another, it is crucial to have a smooth cell handover to sustain continuous connectivity experience.
Beam-level mobility management in 5G relies on periodic channel state information reference signals and synchronization signal blocks for beam sweeping, measurement, and reporting. However, frequent beam sweeping introduces significant latency and high overhead[2], making it impractical for HSR communications. Cell-level mobility management primarily relies on an event A3-based handover process, encompassing handover trigger, preparation, and execution. Before handover completion, the UE continues to communicate with the source base station (BS), often resulting in a sharp decline in the communication quality. Additionally, with a fixed handover period, the link quality between the high-speed terminal and source BS deteriorates faster, increasing the risk of handover failure. Hence, innovative handover mechanisms are urgently needed to address these limitations.
@@
In recent years, deep neural networks (DNNs) have proven their effectiveness in capturing the complex and nonlinear characteristics of wireless channels to enhance mobility management. For example, in [2] the optimal beam for future instances was predicted by using a multi-layer long short term memory (LSTM) network. The appropriate BS for cell handover was predicted by using convolutional neural network (CNN)-based auto-encoders and LSTM networks in [3]. However, these methods either measure all beams, resulting in high overhead, or measure a part of all beams, resulting in notable performance loss. Additionally, existing AI-based cell-level mobility management methods focus solely on the prediction accuracy of the reference signal received power (RSRP), ignoring the communication quality degradation resulting from the reactive nature of the event A3-based handover mechanism. Recently, the 3GPP 5G-Advanced New Radio (NR) standardization has launched studies on AI-based mobility management for highly reliable handovers in Release 19.111https://www.3gpp.org/ftp/tsg_ran/TSG_RAN/TSGR_102/Docs/RP-234055.zip Mobility management methods for HSR communications are expected to leverage case-specific characteristics in the design phase or, alternatively, learn from data to yield improved performance in a data-driven manner.
In this article, we provide two AI-based beam-level and cell-level mobility management approaches for HSR communications. In Section II, we reveal the potential of AI-based spatial-temporal beam prediction for HSR communications. Particularly, by leveraging the sparsity and time-domain correlation of channels in HSR communications, we propose compressed spatial multi-beam measurements via compressive sensing (CS) to improve the spatial-temporal beam prediction accuracy in HSR communications. In Section III, we investigate AI-based cell-level mobility management for HSR scenarios and propose a novel AI-assisted cell handover mechanism, performing handover proactively before the signal quality deteriorates by AI-based handover event prediction. Section V introduces the challenges and opportunities encountered in AI-based mobility management for HSR communications, followed by our conclusions in Section VI. A summary of our overview of AI-based mobility enhancement schemes is given in Table I.
II AI-Based Compressed Multi-Beam Measurement for HSR Communications
II-A AI-Based Beam-Level Mobility Management in HSR Communications
When a user equipment (UE) performs a beam switch within the same cell, it initiates beam-level mobility. This process, which includes beam sweeping, beam measurement, beam reporting, and beam determination is managed within the cell, where the physical (PHY) and media access control (MAC) layers are sufficient to handle it, thereby avoiding the need for radio resource control (RRC) operations. The detailed beam-level mobility management process is shown in Fig. 2.
@@ II-A AI-Based Beam-Level Mobility Management in HSR Communications
Beam-level mobility management enhancement involves three types of beam search methods: traditional methods including exhaustive search and hierarchical search, advanced numerical methods, and AI-based methods. The exhaustive beam search method examines all the candidate beams in the predefined codebook one by one to obtain beam quality information. The hierarchical beam search method provides a more efficient alternative, which first searches within a wide beam range and then further searches for a narrow beam within the angular range of the selected wide beam. However, traditional beam search algorithms suffer from a large measurement overhead, which is not suitable for HSR communications which require frequent beam switching. In addition, enhancement techniques, based on numerical methods have been proposed. More specifically, [4] exploited the tree search method for efficient beam training based on the hierarchical codebook, whilst [5] proposed an optimal beam training model based on Markov’s decision process. These numerical approaches mainly target general communication environments, making it challenging to tailor them for particular communication scenarios. In HSR communications, with some coarse-grained beam quality measurements of the serving cell, AI can be utilized to estimate the finely-grained optimal beam pair in the current instance or predict the optimal beam pair in future instances, which leads to reduced measurement overhead and enhanced throughput.
There are two primary use cases for AI models applied in beam-level mobility management, i.e., temporal beam prediction and spatial beam prediction. For temporal beam prediction, where the beam quality information of previous instances is utilized to predict the quality of beams in one or more future instances, most AI-based methods employ recurrent neural networks (RNNs) and their variants, such as LSTM and gated recurrent unit (GRU) models. This is driven by the ability of RNN models to identify temporal patterns in time series data, allowing them to gather extensive information. A two-layer convolutional LSTM network can be used for predicting the beam quality across spatial and temporal dimensions as detailed in [6]. The combination of a CNN and LSTM network was proposed in [7] for predicting the optimal beam ID across temporal dimensions. A multi-layer LSTM network was applied in [8] for forecasting the optimal future beam. Spatial beam prediction exploits spatial correlations among adjacent beams to estimate the quality of all beams using incomplete beam data. CNNs and networks with fully connected (FC) layers have been commonly employed for this purpose [9, 10, 11, 12]. Due to the high mobility in HSR communications, the signaling overhead and the time required for frequent beam sweeping become prohibitive. For these reasons, temporal or spatial-temporal beam prediction offers greater benefits in terms of resource conservation and stable performance of HSR communications.
Different tasks in beam-level mobility management require distinct inputs to the AI models in HSR communications. In order to instigate further discussion, there are two vital concepts for beam-level mobility management, namely Set A and Set B.222https://www.3gpp.org/ftp/tsg_ran/WG1_RL1/TSGR1_109-e/Inbox/drafts/9.2.3.2/R1-220xxxx%20%5B109-e-R18-AIML-06%5D-Summary-1-v022-MediaTek_Intel.docx Set A comprises all narrow beams from the predefined codebook, while Set B encompasses the beams that have been measured and are utilized as inputs into the AI models. Various alternative model input methods exist with respect to the relationship between Set A and Set B. For example, the AI model could receive the input of RSRP measurements of partial narrow beams, making Set B a subset of Set A. This is the most commonly used method in HSR communications, which greatly reduces the measurement overhead. Another option is to utilize RSRP measurements of wide beams to estimate the quality of narrow beams, where Set B and Set A are distinct. In addition, whenever the RSRP measurements of all narrow beams are the inputs into the AI models, Set B can be the same as Set A. The first two types of inputs, where the inputs to the model are RSRP measurements of some narrow beams and all wide beams, are designed for temporal and spatial beam prediction models, aiming to reduce the measurement overhead by lowering the frequency of beam measurements, which makes them more suitable for latency-sensitive applications. In contrast, the last type of input relies on measuring the RSRP of all candidate beams to conduct temporal beam prediction, achieving the higher model prediction accuracy. This method is well-suited for applications that demand exceptionally high accuracy in prediction models and are less sensitive to latency. Instead of using RSRP, supplementary auxiliary information, like the position of the UE, its mobility vector, the signal-to-noise ratio (SNR), and the beam index can currently be utilized to provide broader spatial context to forecast the state of the link connection [13].
II-B AI-Based Compressed Multi-Beam Measurement Scheme
@@
Here, we investigate DL-based spatial-temporal beam prediction using compressed multi-beam channel measurements. DL is employed to output the beam quality in the future slots using historical multi-beam channel measurements in the compressed domain, where the CS measurement matrix can be learned jointly. In the training phase, the complete beam channel information from historical instances serves as input into the model. The CS matrix serves as the first layer, namely the linear compression layer. The compressed channel information is fed into the remaining neural network to reconstruct the complete RSRP information of all candidate beams for a future instance. In the remaining module, a sub-pixel convolution layer and a 2D upsampling layer are applied before data is fed into the convolutional LSTM (Conv-LSTM) layers. These two layers learn to upsample low-resolution features from limited beam measurements, mapping them to high-resolution feature maps of the complete beam measurements. Then, the initial estimates of the complete beam measurement map are fed into the Conv-LSTM layers to calculate the temporal variations by leveraging historical data and local correlations. Finally, the network outputs a complete RSRP map for an upcoming time slot. As a learnable parameter, the CS measurement matrix is jointly optimized during the training of the model. The structure of the proposed AI model for beam-level mobility management, namely CSAI, is illustrated in Fig. 3(d).
In the inference phase, the trained measurement matrices are applied at the transceivers while the compressed channel measurements are fed into the trained neural network to conduct beam prediction. Meanwhile, the model can undergo periodic retraining or continuous updates through online learning. Detailed deployment information is shown in Fig. 2 Compared with the traditional sweeping method, the multi-beam compression offers to the beam prediction network richer information with the same number of beam measurements.
@@
Now, we compare the performance of AI-based spatial-temporal beam prediction methods, whose structures are shown in Fig. 3. The AI-based methods in Fig. 3(a) [8] and Fig. 3(b) [9] employ the traditional LSTM and CNN architecture, respectively, to predict the high-resolution beams. Compared with the CSAI in Fig. 3(d), the AI-based method in Fig. 3(c) [6] has the same network structure except for the beam measurement module. In the experiments, the channel used in our experiments is generated by QuaDRiGa, and we consider the UMa scenario for outdoor 30 GHz mmWave communications, while the UE follows a straight-line trajectory which aligns with the HSR scenario. Four distinct datasets, representing UE speeds of 60km/h, 120km/h, 350km/h and 500km/h, are utilized in the evaluation. These datasets comprise 40,000 training samples, 5,000 validation samples, and 5,000 test samples. The batch size is set to 256, and the training runs for a maximum of 800 epochs. The learning rate changes dynamically from to , utilizing the adaptive moment estimation optimizer. We consider the prediction accuracy for the best beam with a compression ratio333The compression ratio is defined as the ratio of the number of input beam measurements to the number of all beam measurements. of . The simple non-AI approach straightforwardly uses the best one out of the partially measured present beams as the prediction. For three traditional AI-based beam prediction methods shown in Fig. 3(a) (b) (c), the model’s input uses equal interval sampling from predefined codebook to obtain beam measurement information with a compression ratio of . As shown in Fig. 4, the AI-based spatial-temporal beam prediction methods outperform significantly the traditional non-AI method. As the speed of the movement increases, all methods generally show an anticipated degradation in the prediction accuracy. The CSAI demonstrates superior performance for all the cases. At the speed of 500km/h, the CSAI is still able to yield a prediction accuracy of 83.76%. The Conv-LSTM network shown in Fig. 3(c) requires 3.84M flops and 0.53M parameters, whereas the proposed CSAI model utilizes 3.88M flops and 0.136M parameters. This demonstrates that the proposed model improves prediction performance with a negligible increase in computational cost compared to the fixed down-sampling beam pattern approach.
III AI-Based Proactive Handover Mechanism for HSR Communications
III-A AI-Based Cell-Level Mobility Management in HSR Communications